the larger of two numbers is 4 less than twice the smaller. three times the largerexceeds four times the smaller by 6. find the numbers.
Answer by robertb(5830)

L = 2s-4,
3L = 4s + 6. By substituting the top eqaution into the bottom equation,
3(2s-4) = 4s+6,
6s-12 = 4s+6,
2s = 18,
s = 9, the smaller number.
L = 2*9-4 = 14, the larger number.